# Constacyclic codes with best-known parameters

### v1

@@ 3 Some general constructions for constacyclic codes
where $N_{m}$ is given in Theorem 3.1,
$\Lambda_{m}=\left\lceil\frac{q|S_{m}|}{q-1}\right\rceil-1,$
and
@@
and if moreover $r\leq 2$, then for the second family of constacyclic codes in Theorem 3.1, the lower bound on their minimum distances can be
@@
where
$\Lambda_{m}^{\prime}=\left\lceil\frac{q|S_{m}^{\prime}|}{q-1}\right\rceil-1,$
and the set $S_{m}^{\prime}$ is obtained from $S_{m}$ by replacing $\lceil N_{m}/2\rceil$ with $\lfloor N_{m}/2\rfloor$.
@@
By construction and by the first part of Lemma 2.13, we know that when $N_{1}\geq 2$, all the cyclotomic cosets of $Z_{n,r}$ with coset leader at most $\delta_{\lceil N_{m}/2\rceil}^{(m)}$ are in $T$. When $N_{1}=1$, clearly this fact also holds.
@@
$\delta^{(m)}_{\lceil N_{m}/2\rceil}$ $\geq\frac{nr(1+(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2)r)}{\gcd(nr,q^{m}-1)}$
$=1+(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2)r.$
Notice that the above lower bound is achieved only when all the elements not divisible by $q$ and at most $1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2\right)r$ in $Z_{n,r}$ are coset leaders (because in proving Lemma 2.6, we apply Lemma 2.5). However, this does not always hold. By Lemma 2.8, we know that there are at least $|S_{m}|$ such elements which are not coset leaders. Hence, combining this with Lemma 2.5, we obtain
$\begin{split}\delta^{(m)}_{\lceil N_{m}/2\rceil}&\geq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-1\right)r+(\lceil q|S_{m}|/(q-1)\rceil-2)r\\ &\geq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil+\lceil q|S_{m}|/(q-1)\rceil-3\right)r,\end{split}$ (3.1)
where the set $S_{m}$ has been defined in the theorem.
@@
Finally, when $r\leq 2$, using the second part of Lemma 2.13 and applying similar arguments as the above, we get the desired lower bound for the minimum distances of the second family of constacyclic codes in Theorem 3.1. ∎
@@
$|S_{m}|\geq\sum_{i=(m+1)/2}^{e(m)-1}\sum_{t=1}^{q-1}\left\lfloor\frac{q^{i}-t(q^{m-i}-1)-1}{r(q^{m-i+1}-q)}\right\rfloor,$
@@
7 7 3 1 22876 23232 27299
7 7 3 2 11438 11616 13651
4 Constacyclic codes with prime length
Recall that $q$ is a prime power, $n$ is a positive integer with $\gcd(n,q)=1$ and $r\mid q-1$. In this section, $n$ is a prime.
Theorem 4.1.
Assume that $n$ is a prime with $n\nmid q-1$, and let $m=\mathrm{ord}_{n}(q)$. Then, an infinite family of constacyclic $n,n-1-m\lceil\frac{n-1}{2m}\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2m)\rceil}{q-1}\right\rceil+\Lambda_{m}]$ codes is constructed. Besides, if $2m(q-1)\geq qr$, an infinite family of constacyclic $n,n-m\lfloor\frac{n-1}{2m}\rfloor,\geq\lfloor\frac{q(n-1)}{2m(q-1)}\rfloor]$ codes is constructed.
Proof.
Since $n$ is a prime and $n\nmid q-1$, we have $\gcd(n,q-1)=1$, and so $\gcd(n,r)=1$ (noticing $r\mid q-1$), which implies
$\mathrm{ord}_{nr}(q)=\mathrm{ord}_{n}(q)\mathrm{ord}_{r}(q)=\mathrm{ord}_{n}(q).$
Let $m=\mathrm{ord}_{n}(q)$. Then, combining Lemma 2.1 with Lemma 2.4 and noticing that $n$ is a prime, we obtain that
$\{l_{i}:i\in Z_{n,r}\}=\{1,m\},$
and
$N_{1}=\gcd(n,(q-1)/r)=1\qquad\textrm{and}\qquad N_{m}=\frac{n-1}{m}.$
Hence, the desired results follow directly from Theorems 3.1 and $3.$ ∎
We remark that in Theorem 4.1, if furthermore $r\leq 2$, then by Theorem 3.3, we know that for the second family of constacyclic codes, the lower bound $\lfloor\frac{q(n-1)}{2m(q-1)}\rfloor$ can be replaced by
$\left\lceil\frac{q\lfloor(n-1)/(2m)\rfloor}{q-1}\right\rceil+\Lambda_{m}^{\prime},$
where $\Lambda_{m}^{\prime}$ has been defined in Theorem 3.3.
We also remark that in Theorem 4.1, when (n1)/m(n-1)/m is even, the two families of constacyclic codes have dimensions (n1)/2(n-1)/2 and $(n+1)/2$ respectively.
The lower bounds on minimum distances in Theorem 4.1 depend on the quotient $\frac{n-1}{\mathrm{ord}_{n}(q)}$ when $q$ is fixed. This quotient is from 1 to n1n-1. Particularly, when $q$ is a prime, it is conjectured that there are infinitely many primes $p$ such that $\frac{q^{p}-1}{q-1}$ is also a prime (called the generalized repunit prime). For this case, we put $n=\frac{q^{p}-1}{q-1}$, and then $\mathrm{ord}_{n}(q)=p$, which implies that the lower bounds are of the form $cn/\log_{q}n$ for some positive constant $c$.
In Table LABEL:tab:nprimec, we present some codes constructed in Theorem 4.1 with prime length $n$ and having optimal or best-known parameters. In this table, $m=\mathrm{ord}_{n}(q)$. We emphasize that those codes marked with “*” in the parameters are constructed by using the ceiling function (that is, from the first family of constacyclic codes in the theorem), and the others are constructed by using the floor function (that is, from the second family of constacyclic codes in the theorem). In addition, the column “Lower bound” records the best lower bound in this paper for the minimum distance of the relevant code (here, actually from Theorems 3.3 and 4.1). We follow these rules in all the other tables of constacyclic codes in this paper.
Table 4.1: Codes in Theorem 4.1
$q$ $m$ $n$ $r$ Lower Bose Parameter Optimality
bound distance
22 2020 $41$ 11 22 44 $41,20,10]^{*}$ best-known
22 2020 $41$ 11 22 44 $[41,21,9]$ best-known
22 $23$ $47$ 11 22 55 $47,23,12]^{*}$ best-known
22 $23$ $47$ 11 22 55 $[47,24,11]$ best-known
33 33 $13$ 11 33 44 $13,6,6]^{*}$ optimal
33 33 $13$ 22 33 44 $13,6,6]^{*}$ optimal
33 $12$ $73$ 11 44 55 $73,36,18]^{*}$ best-known
33 $12$ $73$ 11 44 55 $[73,37,17]$ best-known
33 $12$ $73$ 22 55 55 $73,36,18]^{*}$ best-known
33 $12$ $73$ 22 55 77 $[73,37,17]$ best-known
33 $23$ $47$ 11 22 66 $47,23,15]^{*}$ best-known
33 $23$ $47$ 11 22 55 $[47,24,14]$ best-known
33 $23$ $47$ 22 22 66 $47,23,15]^{*}$ best-known
33 $23$ $47$ 22 22 55 $[47,24,14]$ best-known
55 33 $31$ 11 77 88 $[31,16,10]$ optimal
55 33 $31$ 22 77 88 $[31,16,10]$ optimal
55 33 $31$ 44 77 88 $[31,16,10]$ optimal
55 $14$ $29$ 11 22 55 $29,14,12]^{*}$ optimal
55 $14$ $29$ 11 22 55 $[29,15,11]$ optimal
55 $14$ $29$ 22 22 44 $29,14,12]^{*}$ optimal
55 $14$ $29$ 22 22 44 $[29,15,11]$ optimal
55 $14$ $29$ 44 22 55 $29,14,12]^{*}$ optimal
55 $14$ $29$ 44 22 55 $[29,15,11]$ optimal
@@ 5 Constacyclic codes with $n=\frac{q^{p}-1}{rs}$
Assume that $nr=(q^{p}-1)/s$ for some prime $p$, and $\gcd(r,nr/\gcd(nr,q-1))=1$. Then, an infinite family of constacyclic $n,n-\lceil N_{1}/2\rceil-p\lceil N_{p}/2\rceil,\geq\left\lceil\frac{q\lceil N_{p}/2\rceil}{q-1}\right\rceil+\Lambda_{p}]$ codes is constructed, where
@@
then an infinite family of constacyclic $n,n-\lfloor N_{1}/2\rfloor-p\lfloor N_{p}/2\rfloor,\geq\lfloor\frac{qN_{p}}{2(q-1)}\rfloor]$ codes is constructed.
@@
Since $nr=(q^{p}-1)/s$ and $p$ is a prime, by Lemma 2.1 we have that for any $i\in Z_{n,r}$, $l_{i}=1$ or $p$, where $l_{i}=|C_{i}^{(q,nr)}|$. Combining this with Lemma 2.4 (noticing $\gcd(r,nr/\gcd(nr,q-1))=1$), we have
$N_{l}=\begin{cases}\gcd(n,(q-1)/r),&l=1,\\ \frac{n-\gcd(n,(q-1)/r)}{p},&l=p.\end{cases}$
Then, the desired results follow directly from Theorems 3.1 and $3.$ ∎
We remark that in Theorem 5.1, if furthermore $r\leq 2$, then by Theorem 3.3, we know that for the second family of constacyclic codes, the lower bound $\lfloor\frac{qN_{p}}{2(q-1)}\rfloor$ can be replaced by
$\left\lceil\frac{q\lfloor N_{p}/2\rfloor}{q-1}\right\rceil+\Lambda_{p}^{\prime},$
where $\Lambda_{p}^{\prime}$ has been defined in Theorem 3.3.
In Theorem 5.1, there are some special cases where such constacyclic codes have dimension (n1)/2(n-1)/2, or $n/2$, or $(n+1)/2$. We state some of them as examples.
@@
In Theorem 5.1, if moreover $n$ is even, $p$ is an odd prime and $2r\mid q-1$, then both $N_{1}$ and $N_{p}$ are even, and so the first family of constacyclic codes has dimension $n/2$.
Example 5.3.
In Theorem 5.1, if moreover $n$ is odd and $p$ is an odd prime, then $N_{1}$ is odd and $N_{p}$ is even, and so, the first family of constacyclic codes has dimension
$n-\frac{N_{1}+1}{2}-\frac{pN_{p}}{2}=\frac{n-1}{2},$
and the second family of constacyclic codes has dimension
$n-\frac{N_{1}-1}{2}-\frac{pN_{p}}{2}=\frac{n+1}{2}.$
In Table LABEL:tab:qoddprceilc, we list some good codes constructed in Theorem 5.1. In the table, $d_{\text{best}}$ stands for the maximal minimum distance of all known $q$-ary linear codes with relevant length and dimension, and its value is from [15].
@@
$q$ $p$ $s$ $r$ Lower Bose Parameter Optimality
bound distance
@@
3 3 2 1 3 5 $13,6,6]^{*}$ optimal
3 3 2 1 3 4 $[13,7,4]$ almost-optimal
3 3 1 1 6 6 $26,13,8]^{*}$ best-known
3 5 22 1 2 5 $11,5,6]^{*}$ optimal
3 5 22 1 1 4 $[11,6,5]$ optimal
3 5 11 2 2 4 $11,5,6]^{*}$ optimal
3 5 11 2 1 4 $[11,6,5]$ optimal
@@
4 3 3 1 4 6 $21,10,8]^{*}$ best-known
4 3 1 1 13 14 $[63,32,15]$ $d_{\text{best}}=16$
5 3 1 1 25 32 $124,62,32]^{*}$ $d_{\text{best}}=33$
5 5 142 1 3 5 $22,11,8]^{*}$ best-known
5 5 71 2 3 5 $22,11,8]^{*}$ best-known
@@ 6 Constacyclic codes with $n=\Phi_{p^{b}}(q)$
Later on, we will see that in this case the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$, always has exactly two elements. So, we can use Theorem 3.1 to construct constacyclic codes having good lower bounds for their minimum distances.
@@
Assume $p\nmid q-1$ and $n=(q^{p^{b}}-1)/(q^{p^{b-1}}-1)$. Then, an infinite family of constacyclic $n,n-1-p^{b}\lceil(n-1)/(2p^{b})\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p^{b}(q-1)\geq qr$, an infinite family of constacyclic $n,n-p^{b}\lfloor(n-1)/(2p^{b})\rfloor,\geq\lfloor\frac{q(n-1)}{2p^{b}(q-1)}\rfloor]$ codes is constructed.
@@
First, we want to determine $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. Since $r\mid q-1$ and $b\geq 1$, we write
$nr=\frac{q^{p^{b}}-1}{(q^{p^{b-1}}-1)/r},$ (6.1)
where the denomenator is a positive integer. Combining (6.1) with Lemma 2.1, we have
$l_{i}\mid p^{b}\quad\textrm{for}\quad i\in Z_{n,r}.$
@@
$\gcd(n,q^{p^{j}}-1)=\gcd(p,q^{p^{j}}-1)=1\quad\textrm{for}\quad j=0,1,\ldots,b-1.$
Hence, we get
$\gcd(n,q^{p^{j}}-1)=1\qquad\textrm{for}\qquad j=0,1,\ldots,b-1,$ (6.3)
@@
$\gcd\left(r,\frac{nr}{\gcd(nr,q^{j}-1)}\right)=\gcd\left(r,\frac{n}{\gcd(n,(q^{j}-1)/r)}\right)=1.$
Thus, applying Lemma 2.3 and (6.3) and using the definition of the Möbius function, we obtain
@@
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}$
$=\frac{n-1}{p^{b}}.$
Moreover, noticing $N_{1}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have
$\{l_{i}:\,i\in Z_{n,r}\}=\{1,p^{b}\}.$
Then, the desired result follows directly from Theorems 3.1 and $3.$ ∎
@@
Assume $p\mid(q-1)/r$ and $n=(q^{p^{b}}-1)/(q^{p^{b-1}}-1)$. Then, an infinite family of constacyclic $n,n-\lceil p/2\rceil-p^{b}\lceil\frac{n-p}{2p^{b}}\rceil,\geq\left\lceil\frac{q\lceil(n-p)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p^{b}(q-1)\geq qr$, an infinite family of constacyclic $n,n-\lfloor p/2\rfloor-p^{b}\lfloor\frac{n-p}{2p^{b}}\rfloor,\geq\lfloor\frac{q(n-p)}{2p^{b}(q-1)}\rfloor]$ codes is constructed.
@@
First, we want to determine the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. As before, we have $l_{i}\mid p^{b}$ for any $i\in Z_{n,r}$.
@@
In addition, by (6.2) and noticing $r\mid q-1$, we have
$\gcd(n,r)=\gcd(p,r).$
So, $\gcd(n,r)=1$ or $p$. If $\gcd(n,r)=1$, clearly we have $\gcd(n/p,r)=1$. If $\gcd(n,r)=p$, then noticing $p\mid(q-1)/r$, we can write $q=ap^{2}+1$ for some positive integer $a$, and so,
@@
$\begin{split}\gcd(r,\frac{nr}{\gcd(nr,q^{p^{j}}-1)})&=\gcd(r,\frac{n}{\gcd(n,(q^{p^{j}}-1)/r)})\\ &=\gcd(r,n/p)=1.\end{split}$ (6.6)
Now, applying Lemma 2.3 with (6.4) and (6.6) and using the definition of the Möbius function, we obtain
@@
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}$
$=\frac{n-p}{p^{b}}.$
Moreover, noticing $N_{1}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have
$\{l_{i}:\,i\in Z_{n,r}\}=\{1,p^{b}\}.$
Then, the desired result follows directly from Theorems 3.1 and $3.$ ∎
@@
Assume $p\mid q-1$, $p\nmid(q-1)/r$, $b\geq 2$ and $n=(q^{p^{b}}-1)/(q^{p^{b-1}}-1)$. Then, an infinite family of constacyclic $n,n-p-p^{b}\lceil\frac{n-p}{2p^{b}}\rceil,\geq\left\lceil\frac{q\lceil(n-p)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.3. Moreover, if either $r\leq 3$ or $2p^{b-1}(q-1)\geq qr$, an infinite family of constacyclic $n,n-p^{b}\lfloor\frac{n-p}{2p^{b}}\rfloor,\geq\lfloor\frac{q(n-p)}{2p^{b}(q-1)}\rfloor]$ codes is constructed.
@@
First, we want to determine the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}^{(q,nr)}|$. As before, we have $l_{i}\mid p^{b}$ for any $i\in Z_{n,r}$.
Since $p\mid q-1$ and $p\nmid(q-1)/r$, we must have $p\mid r$.
@@
where the last congruence follows from the assumption $b\geq 2$. So, we have $p\nmid n/p$. Notice that $\gcd(n,r)=\gcd(p,r)=1$ or $p$. Then, we must have
@@
Then, using (6.7) and noticing $p\mid r$, we obtain
$\begin{split}\gcd\left(r,\frac{nr}{\gcd(nr,q-1)}\right)&=\gcd\left(r,\frac{n}{\gcd(n,(q-1)/r)}\right)\\ &=\gcd(r,n)=\gcd(r,p)=p.\end{split}$ (6.10)
@@
Now, applying Lemma 2.3 with (6.8), (6.10) and (6.11) and using the definition of the Möbius function, we obtain

### v2

@@ 3 Some general constructions for constacyclic codes
where $N_{m}$ is given in Theorem 3.1, $\Lambda_{m}=\left\lceil\frac{q|S_{m}|}{q-1}\right\rceil-1,$ and
@@
and in addition, if $l=1$ and $r\leq 2$, for the second family of $q$-ary constacyclic codes in Theorem 3.1, the lower bound on their minimum distances can be
@@
where $\Lambda_{m}^{\prime}=\left\lceil\frac{q|S_{m}^{\prime}|}{q-1}\right\rceil-1,$ and the set $S_{m}^{\prime}$ is obtained from $S_{m}$ by replacing $\lceil N_{m}/2\rceil$ with $\lfloor N_{m}/2\rfloor$.
@@
By construction and by the first part of Lemma 2.13, we know that when $l=1$ and $N_{1}\geq 2$, all the cyclotomic cosets of $Z_{n,r}$ with coset leader at most $\delta_{\lceil N_{m}/2\rceil}^{(m)}$ are in $T$. When $N_{l}=1$, clearly this fact also holds.
@@
$\delta^{(m)}_{\lceil N_{m}/2\rceil}\geq\frac{nr(1+(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2)r)}{\gcd(nr,q^{m}-1)}=1+(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2)r.$
Notice that the above lower bound is achieved only when all the elements not divisible by $q$ and at most $1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-2\right)r$ in $Z_{n,r}$ are coset leaders (because in proving Lemma 2.6, we apply Lemma 2.5). However, this does not always hold. By Lemma 2.8, we know that there are at least $|S_{m}|$ such elements which are not coset leaders, where the set $S_{m}$ has been defined in the theorem. Hence, combining this with Lemma 2.5, we obtain
$\begin{split}\delta^{(m)}_{\lceil N_{m}/2\rceil}&\geq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil-1\right)r+(\lceil q|S_{m}|/(q-1)\rceil-2)r\\ &\geq 1+\left(\lceil q\lceil N_{m}/2\rceil/(q-1)\rceil+\lceil q|S_{m}|/(q-1)\rceil-3\right)r.\end{split}$ (3.1)
@@
Finally, when $l=1$ and $r\leq 2$, using the second part of Lemma 2.13 and applying similar arguments as the above, we get the desired lower bound for the minimum distances of the second family of constacyclic codes in Theorem 3.1. ∎
@@
$|S_{m}|\geq\sum_{i=(m+1)/2}^{e(m)-1}\sum_{j=1}^{q-1}\left\lfloor\frac{q^{i}-j(q^{m-i}-1)-1}{r(q^{m-i+1}-q)}\right\rfloor,$
@@ 4 Constacyclic codes with $n=\frac{q^{p}-1}{rs}$
Assume that $nr=(q^{p}-1)/s$ for some prime $p$, and $\gcd(r,nr/\gcd(nr,q-1))=1$. Then, an infinite family of $q$-ary $n,n-\lceil N_{1}/2\rceil-p\lceil N_{p}/2\rceil,\geq\left\lceil\frac{q\lceil N_{p}/2\rceil}{q-1}\right\rceil+\Lambda_{p}]$ constacyclic codes is constructed, where
@@
then an infinite family of $q$-ary $n,n-\lfloor N_{1}/2\rfloor-p\lfloor N_{p}/2\rfloor,\geq\lfloor\frac{qN_{p}}{2(q-1)}\rfloor]$ constacyclic codes is constructed.
@@
Since $nr=(q^{p}-1)/s$ and $p$ is a prime, by Lemma 2.1 we have that for any $i\in Z_{n,r}$, $l_{i}=1$ or $p$, where $l_{i}=|C_{i}|$. Combining this with Lemma 2.4 (noticing $\gcd(r,nr/\gcd(nr,q-1))=1$), we have
$N_{1}=\gcd(n,(q-1)/r),\quad N_{p}=(n-\gcd(n,(q-1)/r))/p.$
Then, the desired results follow directly from Theorems 3.1 and 3.4. ∎
We remark that in Theorem 4.1, since $N_{1}>0$, by Remark 3.2 we know that the constacyclic codes in Theorem 4.1 are monomially equivalent to some cyclic codes over $\mathbb{F}_{q}$.
We notice that some codes in Theorem 4.1 are better than BCH constacyclic codes; see Example 4.2.
@@
Let $q=7$, $p=3$, $s=9$, $r=1$, and $n=38$. Then, the $7$-cyclotomic cosets in $Z_{38,1}$ are: $C_{1}=\{1,7,11\}$, $C_{2}=\{2,14,22\}$, $C_{3}=\{3,21,33\}$, $C_{4}=\{4,6,28\}$, $C_{5}=\{5,17,35\}$, $C_{8}=\{8,12,18\}$, $C_{9}=\{9,23,25\}$, $C_{10}=\{10,32,34\}$, $C_{13}=\{13,15,29\}$, $C_{16}=\{16,24,36\}$, $C_{19}=\{19\}$, $C_{20}=\{20,26,30\}$, $C_{27}=\{27,31,37\}$, $C_{38}=\{38\}$. So, $\{l_{i},i\in Z_{38,1}\}=\{1,3\}$, $N_{1}=2$ and $N_{3}=12$. From Theorem 4.1, we take the defining set as the set $C_{1}\cup C_{2}\cup C_{3}\cup C_{4}\cup C_{5}\cup C_{8}\cup C_{19}$ and obtain a $7$-ary $[38,19,13]$ cyclic code, which is not a BCH code and has the best-known parameters [18]. In this case, there are several 7-ary BCH codes of length 38 and dimension 19, but their minimum distances are at most 12.
In Table LABEL:tab:qp, we list some good codes constructed in Theorem 4.1. In the table, those codes marked with “*” in the parameters are constructed by using the ceiling function (that is, from the first family of constacyclic codes in the theorem), and the others are constructed by using the floor function (that is, from the second family of constacyclic codes in the theorem). In addition, the column “Lower bound” records the best lower bound in this paper for the minimum distance of the relevant code (here, actually from Theorems 3.1, 3.4 and 4.1). We follow these rules in all the other tables of constacyclic codes in this paper.
@@
$q$ $p$ $s$ $r$ Lower bound Bose distance Parameter Optimality
@@ 5 Constacyclic codes with $n=\Phi_{p^{b}}(q)$
Later on, we will see that in this case the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}|$, always has exactly two elements. So, we can use Theorem 3.1 to construct constacyclic codes having good lower bounds for their minimum distances.
@@
Assume $p\nmid q-1$ and $n=\Phi_{p^{b}}(q)$. Then, an infinite family of $q$-ary $n,n-1-p^{b}\lceil(n-1)/(2p^{b})\rceil,\geq\left\lceil\frac{q\lceil(n-1)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ constacyclic codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.4. Moreover, if either $r\leq 3$ or $2p^{b}(q-1)\geq qr$, an infinite family of $q$-ary $n,n-p^{b}\lfloor(n-1)/(2p^{b})\rfloor,\geq\lfloor\frac{q(n-1)}{2p^{b}(q-1)}\rfloor]$ constacyclic codes is constructed.
@@
First, we want to determine the set $\{l_{i}:\,i\in Z_{n,r}\}$, where $l_{i}=|C_{i}|$. Since $n=\Phi_{p^{b}}(q)$ and $r\mid q-1$, by Lemma 2.1 we have $l_{i}\mid p^{b}$ for any $i\in Z_{n,r}$.
@@
$\gcd(n,q^{p^{j}}-1)=\gcd(p,q^{p^{j}}-1)=1\quad\textrm{for}\quad j=0,1,\ldots,b-1,$ (5.2)
@@
$\gcd\left(r,\frac{nr}{\gcd(nr,q^{p^{j}}-1)}\right)=\gcd\left(r,\frac{n}{\gcd(n,(q^{p^{j}}-1)/r)}\right)=1.$ (5.3)
Thus, applying Lemma 2.3 with (5.2) and (5.3) and noticing the definition of the Möbius function, we obtain
@@
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}=\frac{n-1}{p^{b}}.$
Moreover, noticing $N_{1}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have $\{l_{i}:\,i\in Z_{n,r}\}=\{1,p^{b}\}.$ Then, the desired result follows directly from Theorems 3.1 and 3.4. ∎
@@
Assume $p\mid(q-1)/r$ and $n=\Phi_{p^{b}}(q)$. Then, an infinite family of $q$-ary $n,n-\lceil p/2\rceil-p^{b}\lceil\frac{n-p}{2p^{b}}\rceil,\geq\left\lceil\frac{q\lceil(n-p)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ constacyclic codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.4. Moreover, if either $r\leq 3$ or $2p^{b}(q-1)\geq qr$, an infinite family of $q$-ary $n,n-\lfloor p/2\rfloor-p^{b}\lfloor\frac{n-p}{2p^{b}}\rfloor,\geq\lfloor\frac{q(n-p)}{2p^{b}(q-1)}\rfloor]$ constacyclic codes is constructed.
@@
In addition, by (5.1) and noticing $r\mid q-1$, we have $\gcd(n,r)=\gcd(p,r).$ So, $\gcd(n,r)=1$ or $p$. If $\gcd(n,r)=1$, clearly we have $\gcd(n/p,r)=1$. If $\gcd(n,r)=p$, then noticing $p\mid(q-1)/r$, we can write $q=ap^{2}+1$ for some positive integer $a$, and so,
@@
$\begin{split}\gcd(r,\frac{nr}{\gcd(nr,q^{p^{j}}-1)})=\gcd(r,\frac{n}{\gcd(n,(q^{p^{j}}-1)/r)})=\gcd(r,n/p)=1.\end{split}$ (5.6)
Now, applying Lemma 2.3 with (5.4) and (5.6), we obtain
@@
$=\frac{\gcd(n,(q^{p^{b}}-1)/r)-\gcd(n,(q^{p^{b-1}}-1)/r)}{p^{b}}=\frac{n-p}{p^{b}}.$
Moreover, noticing $N_{1}+p^{b}N_{p^{b}}=n=|Z_{n,r}|$, we must have $\{l_{i}:\,i\in Z_{n,r}\}=\{1,p^{b}\}.$ Then, the desired result follows directly from Theorems 3.1 and 3.4. ∎
We remark that in Theorems 5.1 and 5.2, since $N_{1}>0$, by Remark 3.2 we know that the constacyclic codes there are monomially equivalent to some cyclic codes over $\mathbb{F}_{q}$.
@@
Assume $p\mid q-1$, $p\nmid(q-1)/r$, and $n=\Phi_{p^{b}}(q)$. Assume further either $b\geq 2$ or $p$ is odd. Then, an infinite family of $q$-ary $n,n-p-p^{b}\lceil\frac{n-p}{2p^{b}}\rceil,\geq\left\lceil\frac{q\lceil(n-p)/(2p^{b})\rceil}{q-1}\right\rceil+\Lambda_{p^{b}}]$ constacyclic codes is constructed, where $\Lambda_{p^{b}}$ has been defined in Theorem 3.4. Moreover, if either $r\leq 3$ or $2p^{b-1}(q-1)\geq qr$, an infinite family of $q$-ary $n,n-p^{b}\lfloor\frac{n-p}{2p^{b}}\rfloor,\geq\lfloor\frac{q(n-p)}{2p^{b}(q-1)}\rfloor]$ constacyclic codes is constructed.
@@
First, since $p\mid q-1$ and $p\nmid(q-1)/r$, we must have $p\mid r$.
@@
where the last congruence follows from the assumption that $b\geq 2$ or $p$ is odd. So, we have $p\nmid n/p$. In addition, using (5.1) and noticing $p\mid r$, we get
$\gcd(n,r)=\gcd(p,r)=p.$ (5.9)
Hence, we must have
@@
Then, using (5.7) and (5.9), we obtain
$\begin{split}\gcd\left(r,\frac{nr}{\gcd(nr,q-1)}\right)&=\gcd\left(r,\frac{n}{\gcd(n,(q-1)/r)}\right)\\ &=\gcd(r,n)=p.\end{split}$ (5.11)
@@
Now, applying Lemma 2.3 with (5.8), (5.11) and (5.12), we obtain